OpenStudy (anonymous):

differentiate y=sqrt(3x^2-cos^2 4x

OpenStudy (anonymous):

write it out as y= (3x^2 - cos^2 4x)^(1/2) so you can use the power rule/chain rule...

OpenStudy (anonymous):

\[\large y=(3x^2-\cos^24x)^{\frac {1}{2}}\] \[\large y^'=\frac {1}{2}(3x^2-\cos^24x)^{-\frac {1}{2}}[3x^2-\cos^24x]^'\]

OpenStudy (anonymous):

my latex is too slow to take the derivative of that part in the brackets... :( but can you do it from here?

OpenStudy (anonymous):

no i cannt

OpenStudy (anonymous):

ok.. we'll just work with the derivative of what's in the brackets then.. ok?

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

[ 3x^2−cos^2 4x ]′ can you do anything with this derivative?